OpenStudy (anonymous):

y=-2/3x+5 x being -2

OpenStudy (anonymous):

can someone please solve the following problem please? y=-2/3x+5 x being -2

OpenStudy (m):

y = 2

OpenStudy (anonymous):

can you show how you got your answer please

OpenStudy (m):

just plug in x with -2 y = -2/(3(-2)+5)

OpenStudy (m):

=-2/(-6+5) = -2/-1 = 2 minus divided by minus is positive